Two forces are acting at an angle 120 deg. The bigger force is 40N and the resultant is perpendicular to the smaller one. The smaller force is

Grade:12th pass

2 Answers

Arun
6 years ago
Dear Rohit
 
Answer is 20 N
 
Solution-
Tan (α) = psin (θ)/p + qcos (θ).
Alpha = 90, tan90 = 1/0.
θ = 120 = 180 - 120 = 60, sin60 = root3/2.
Cos60 = 1/2.
1/0 = p (root3/2)/p-40 (1/2).
P = 20.

Because cos (180-θ) = -cos (θ).
 
 
Regards
Arun (askIITians forum expert)
Kushagra Madhukar
3 years ago
Dear student,
Please find the attached solution to your problem.
 
tan (α) = psin(θ) / {p + qcos(θ)}.
α = 90, tan90 = 1/0.
θ = 120 = 180 - 120 = 60, sin60 = 3/2.
Cos60 = – 1/2.
1/0 = p (3/2) / {p – 40(1/2)}.
p = 20.
 
Hope it helps.
Thanks and regards,
Kushagra